OpenStudy (anonymous):

f(x)=5+7x^2 Does the f(x) represent a polynomial function?

Parth (parthkohli):

What do you think?

OpenStudy (anonymous):

Yes

Parth (parthkohli):

How so?

OpenStudy (anonymous):

because of the equation

Parth (parthkohli):

I mean, why do you think it is a polynomial function?

OpenStudy (anonymous):

because of the terms? I really honestly don't know if it is

OpenStudy (anonymous):

If it is a polynomial function what is it's type?

Parth (parthkohli):

Well, does it look like this thing?\[ax^2 + bx + c\]If it does, then it's a polynomial function. Remember that \(a,b,c\) can hold the value \(0\) too.

OpenStudy (anonymous):

It somewhat looks like that, except that it's flipped

Parth (parthkohli):

Yeah, but it still is. So it's a polynomial :-)

OpenStudy (anonymous):

Okay so what wold the polynomials type be? linear, cubic, quadratic?

OpenStudy (anonymous):

A polynomial function is defined as a function f(x) = ax^n + bx^(n-1) + ... + cx^(n-(n-2)) + dx^(n-(n-1)) + ex^(n-n) = ax^n + bx^(n-1) + ... +cx^2 + dx + e. In your equation, c=7 and e=5 are coefficients, and the rest of the coefficients are equal to zero so those terms don't appear. Therefore your f(x) represents a polynomial function.

Parth (parthkohli):

What is the highest power of \(x\) which you see?

Parth (parthkohli):

If the highest power is 1, then linear. If 2, then quadratic. If 3, then cubic. If 4, then quartic.

OpenStudy (anonymous):

do you mean poly?

OpenStudy (anonymous):

Oh, thank you all you really made me understand this a lot better

OpenStudy (anonymous):

If it fits the definition of a polynomial function, then it can be called a polynomial function. Math is mostly a lot of definitions! :)